OpenStudy (anonymous):

solve each system of equations. 5/x^2-2/y^2+3=0 3/x^2+1/y^2=7

jhonyy9 (jhonyy9):

5y2 -2x2 +3x2y2 =0 3y2 +x2 -7x2y2 =0 -2x2 +3x2y2 +5y2 =0 x2 -7x2y2 +3y2 =0